# Majority voting is not good for heaven or hell, with mirrored performance

### v3

@@ 2 Notation and problem formulation
$\ell$ number of individualists
$g$ size of the group
$n=g+\ell$ number of agents
@@
${X}={X}^{\mu}$ gain generator having a mean $\mu$ and a finite variance
@@
${\mathbf{{X}}}={\mathbf{{X}}}^{\mu}$ proposal generator corresponding to ${X}^{\mu}$
@@
${\mathbf{{X}}}=({{X}_{1}},\ldots,{{X}_{n}})$ generator ${\mathbf{{X}}}$ with ${{X}_{1}},\ldots,{{X}_{g}}$ corresponding to the group
@@
$n^{+}({\mathbf{{x}}})$ number of votes cast for proposal ${\mathbf{{x}}}$
$I({\mathbf{{x}}})$ implementation probability of proposal ${\mathbf{{x}}}$
The purpose of this study is to find out how the effectiveness of majority decisions depends on the favorability of the environment. In particular, we examine the relationship between ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}_{1}}\bigr{)}$ and ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}_{2}}\bigr{)}$ for arbitrary opposite proposal generators ${\mathbf{{X}}}_{1}$ and ${\mathbf{{X}}}_{2}$, with two different notions of opposition.
@@ 3.1 Relationship between gains under opposite proposals
Let ${\mathbf{{Y\!}}_{\mathbf{{X}}}}$ be the function of proposal generator ${\mathbf{{X}}}$ defined by
${\mathbf{{Y\!}}_{\mathbf{{X}}}}(\omega)={{\mathbf{{X}}}(\omega)}\hskip 1.1ptI({{\mathbf{{X}}}(\omega)}),\quad\omega\in\Omega.$ (1)
@@
Since the voting body is antisymmetric,
@@
By (1), (2), and (3), for any ${\mathbf{{X}}}$, it holds that
@@
$=$ ${\mathbb{E}}\bigl{(}-{{\mathbf{{X}}}(\omega)}(1-I({{\mathbf{{X}}}(\omega)}))\bigr{)}={\mathbb{E}}(-{{\mathbf{{X}}}(\omega)})+{\mathbb{E}}({{\mathbf{{X}}}(\omega)}\hskip 1.1ptI({{\mathbf{{X}}}(\omega)}))$
$=$ $-{\mathbb{E}}({\mathbf{{X}}})+{\mathbb{E}}\bigl{(}{\mathbf{{Y\!}}_{\mathbf{{X}}}}(\omega)\bigr{)}=-\mu\bm{1}+{\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}}).$
This completes the proof. ∎
@@
Any voting body formed by individualists, members of one group, and the symmetrized majority rule is antisymmetric.
@@
Observe that $\tfrac{1}{2}(\mathop{\rm sgn}\nolimits{{x}_{i}}+1)$ is the number of votes cast by individualist $i$ for proposal ${\mathbf{{x}}}=({{x}_{1}},\ldots,{{x}_{n}}).$ Similarly, $\tfrac{1}{2}\left(\mathop{\rm sgn}\nolimits\sum_{i=1}^{g}{{x}_{i}}+1\right)$ is the number of votes cast by each group member. Then for any voting body of the given type and any ${\mathbf{{x}}}={{\mathbf{{X}}}(\omega)},\,\omega\in\Omega$, we have
$n^{+}({\mathbf{{x}}})+n^{+}(-{\mathbf{{x}}})$ $=$ $\tfrac{1}{2}\left(\mathop{\rm sgn}\nolimits\sum_{i=1}^{g}{{x}_{i}}+1\right)g+\tfrac{1}{2}\!\sum_{i=g+1}^{n}(\mathop{\rm sgn}\nolimits{{x}_{i}}+1)$
$+$ $\tfrac{1}{2}\left(\mathop{\rm sgn}\nolimits\sum_{i=1}^{g}(-{{x}_{i}})+1\right)g+\tfrac{1}{2}\!\sum_{i=g+1}^{n}(\mathop{\rm sgn}\nolimits(-{{x}_{i}})+1)$
$=$ $g+\ell=n.$
Therefore, matching the corresponding relations for $n^{+}({\mathbf{{x}}})$ and $n^{+}(-{\mathbf{{x}}})$ we have
$n^{+}({\mathbf{{x}}})>(=)\,(<)\;\tfrac{n}{2}\,\Leftrightarrow\,n^{+}(-{\mathbf{{x}}})<(=)\,(>)\;\tfrac{n}{2}\,.$ (4)
Since by the definition of symmetrized majority,
$n^{+}({\mathbf{{x}}})>(=)\,(<)\;\tfrac{n}{2}\,\Leftrightarrow\,I({\mathbf{{x}}})=1\,\bigl{(}\tfrac{1}{2}\bigr{)}\,(0),$
it follows from (4) that $I({\mathbf{{x}}})+I(-{\mathbf{{x}}})=1$. Thus, the voting body is antisymmetric. ∎
@@
For any gain generator ${X}$ with mean $\mu$ and for any voting body formed by individualists, members of one group, and the symmetrized majority rule, it holds that
@@
Theorem 3 follows from Lemmas 1 and 2.
Remark 1.
Similarly, it can be shown that Theorem 3 extends to all voting bodies consisting of individualists, any number of groups, and the symmetrized majority rule. Moreover, each group may vote either to increase the total group capital ((((as previously assumed)))) or under the intra-group symmetrized majority rule. In the latter case, every group member supports the decision made by the intra-group symmetrized majority, and the group is called an alternative one (while the term “group” is retained for groups of the former type). Since an individualist is a special case of a group of either type, the formulation reduces to: Theorem 3 extends to all voting bodies formed by any number of groups of arbitrary types and the symmetrized majority rule.
@@ 3.2 Gains on location families of generators
Suppose that society consists of individualists and a group, the social decision rule is symmetrized majority, and the distribution of the gain generator ${X}^{\mu}$ is symmetric about $\mu.$ Then
@@
Like Theorem 3, Theorem 4 can be extended as described in Remark 1.
@@ 4 Interpretation and discussion
For any social decision rule $\psi,$ let the expected relative capital gain (ERCG) of an agent under $\psi$ be the mathematical expectation of the difference between the agent’s capital gain under $\psi$ and that under the basic rule.
For families (5), consider the ERCG as a function of $\mu.$
@@
For any family (5) and any agent, the expected capital gain under the basic rule is 00 when $\mu\leq 0$ and $\mu$ when $\mu>0.$ Therefore, for any $\mu>0,$ the left-hand and right-hand sides of (6) are the vectors of agents’ expected relative capital gains under gain generators ${X}^{-\mu}={X}^{\mu}-2\mu$ and ${X}^{\mu}$, respectively. Since by Theorem 4 they are equal, the ERCG is an even function of $\mu.$
@@
For the society of 21 individualists with Gaussian generators (5), the expected capital gain under simple majority rule was shown in Fig. 1. The ERCG for this family is presented in Fig. 2a. In addition to a pit of losses in unfavorable environments, this function has a mirror image of that pit in favorable environments. Decisions made by majority votes in these two areas are, on average, disadvantageous to society compared to the basic rule.
The reason for the pit of losses paradox is that in unfavorable environments, the average positive gain is smaller, while the average absolute loss is larger than in the neutral environment of the same variance (where these average values are equal). Therefore, the average difference between [the absolute values of] a gain and a loss is negative. Consequently, the proposals accepted by simple majority with a small margin of votes in favor (which is typical in an unfavorable environment) yield, on average, a total loss. Thus, in such environments, the majority’s gain does not compensate the loss suffered by the minority. Note that this imbalance identified using the ViSE model is relevant to practice.
@@
Fig. 2a allows us to conclude that if the set of environments with $\mu/\sigma<-0.254$ is considered as “hell” and the set of environments with $\mu/\sigma>0.254$ as “heaven”, then the simple majority rule is worse than the basic rule in both of them. Simple majority outperforms the basic rule only in a vicinity of the neutral environment ($\mu=0$), where the ERCG peak is quite high.
For the society with a group of 7 agents and 14 individualists (Fig. 2b), Gaussian gain generators produce a completely positive ERCG for group members (blue curve). In contrast, the ERCG of individualists is negative, except for a narrow interval of $\mu/\sigma\in[-0.03,0.03]$ (red curve).
It is instructive to compare this with the ERCG in the societies of 21 individualists (red dotted curve) and of 21 group members (blue dotted curve). The most advantageous position is in the group of 7 agents and the worst is among the individualists complementing this group. Such a group can be considered as an “elite”. It usually receives relatively high capital gains at the expense of individualists. It should be noted that the ERCG of the whole mixed society (gray curve) is even lower than that for the society of 21 individualists. The impact of a “responsible elite” on public welfare was investigated in [43, 44]. In most cases, such an elite provides society with capital growth.
Obviously, the most beneficial strategy for the whole society is to be one cohesive group. However, in practice this is unrealistic. While a group of like-minded people is able to develop a common objective function, this is problematic for society. Therefore, non-consensus-based decisions are of primary interest. The “basic rule”, which outperforms the simple majority in both “hell” and “heaven” is extremely crude. How can taking into account voters’ opinions improve the outcome? Such a social decision rule is voting with the optimal threshold ([9, 26]). This threshold is exactly $0.5$ in the neutral environment, as well as for $\mu/\sigma\in[-0.1,0.1]$ in our example (Fig. 2a). Outside a certain interval around 00, the optimal threshold is usually (but not always) higher in unfavorable and lower in favorable environments. The ERCG curve in the case of voting with the optimal threshold clearly has no pit of losses. For the society of 21 individualists, it is shown in Fig. 2a.
@@ 5 Conclusion
The ViSE model is designed to evaluate the effectiveness of individual voting strategies and social decision rules in environments of varying favorability and, more generally, in environments with various gain distributions. In this paper, we show that in the case of Gaussian environments and two-component (a group and individualists) societies, the symmetrized majority rule is effective only under more or less neutral conditions, while in “heaven” or “hell” it is harmful. Furthermore, for environments with symmetric distributions, the performance of this rule is symmetric about the neutrality point. For general distributions with finite mean, the performance is the same for any two opposite generators.
@@
Acknowledgments We are very grateful to two anonymous referees, especially, for suggesting the idea of reverse time to obtain a simpler proof of the technical result. In the final version, we use a very similar approach in other terms, supplemented by the condition of antisymmetry of the voting body, which is necessary to generalize the result. We also thank the participants of the seminar on stochastic voting models at the Institute of Control Sciences of the Russian Academy of Sciences: Yana Tsodikova, Anton Loginov, Vitaly Malyshev, Vladislav Maksimov, Anna Khmelnitskaya, Irina Kalushina, Zoya Lezina, Vladimir Borzenko.
@@
Funding   Work of P.C. was supported by the European Union (ERC, GENERALIZATION, 101039692). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.

### v4

@@ 2 Basic notation and problem formulation
$\{1,\ldots,n\}$ society; $n\in{\mathbb{N}}$
$\{1,\ldots,g\}$ group; $0\leq g\leq n$
$\{g+1,\ldots,n\}$ set of individualists; $\ell=n-g$ is the number of them
@@
${X}={X}^{\mu}$ gain generator with mean $\mu$
@@
${\mathbf{{X}}}=({{X}_{1}},\ldots,{{X}_{n}})$ proposal generator with gain generators ${{X}_{1}},\ldots,{{X}_{n}}$
${\mathcal{X}}^{\mu}$ set of proposal generators whose components are i.i.d.
gain generators with mean $\mu$
${\mathbf{{X}}}^{\mu}{\color[rgb]{0,0,0}\,\in{\mathcal{X}}^{\mu}}$ proposal generator corresponding to gain generator ${X}^{\mu}$
@@
${\color[rgb]{0,0,0}n_{i}^{+}({\mathbf{{x}}}),\,}n^{+}({\mathbf{{x}}})$ number of votes cast for ${\mathbf{{x}}}$ by agent $i$ and society, resp.
$I({\mathbf{{x}}})$ implementation probability of a fixed proposal ${\mathbf{{x}}}\in{\mathbb{R}}^{n}$
The purpose of this study is to find out how the effectiveness of majority decisions depends on the favorability of the environment. We examine the relationship between ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}_{1}}\bigr{)}$ and ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}_{2}}\bigr{)}$ for opposite proposal generators ${\mathbf{{X}}}_{1}$ and ${\mathbf{{X}}}_{2}$, with two different notions of opposition. Numerical examples correspond to the normal distribution.
@@ 3.1 Relationship between gains under opposite proposals
For an arbitrary given ${\mathbf{{X}}}\in{\mathcal{X}}^{\mu},$ we define ${\mathbf{{Y\!}}_{\mathbf{{X}}}}={\mathbf{{X}}}\hskip 1.1ptI({\mathbf{{X}}})$.
Since the voting body is antisymmetric, it follows from (1) that
${\mathbf{{Y\!}}_{\mathbf{{X}}}}-{\mathbf{{Y\!}}_{-{\mathbf{{X}}}}}={\mathbf{{X}}}\hskip 1.1ptI({\mathbf{{X}}})-(-{\mathbf{{X}}})I(-{\mathbf{{X}}})={\mathbf{{X}}}(I({\mathbf{{X}}})+I(-{\mathbf{{X}}}))={\mathbf{{X}}}.$ (2)
@@
By (3) and (2), it holds that
${\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})-{\mathbb{E}}({\mathbf{D}}_{-{\mathbf{{X}}}})={\mathbb{E}}({\mathbf{{Y\!}}_{\mathbf{{X}}}})-{\mathbb{E}}({\mathbf{{Y\!}}_{-{\mathbf{{X}}}}})={\mathbb{E}}({\mathbf{{Y\!}}_{\mathbf{{X}}}}-{\mathbf{{Y\!}}_{-{\mathbf{{X}}}}})={\mathbb{E}}{\mathbf{{X}}}=\mu\bm{1}.$
Remark 1.
@@
The decision-making process with generator ${\mathbf{{X}}}$, viewed in the reverse order, shares the generator with the forward process with $-{\mathbf{{X}}}$. However, under the antisymmetry of the voting body, these processes are complementary in terms of decisions made. This observation, due to the Reviewer, allows us to state that the key antisymmetry condition is equivalent to the requirement that the forward and corresponding backward transitions are evaluated by the voting body oppositely in the sense of Eq. (1).
One more interesting observation by the Reviewer is that Lemma 1 still holds for another model where proposals are generated in the same way, but agents’ votes on proposals (which are changes to the state of the world) stem from their preferences over the states of the world themselves (cf. Subsection 1.3). This origin of votes (or social decisions) along with appropriate tie-breaking rules determine the reversal symmetry of the agents’ strategies (respectively, of the voting body). When it is satisfied, state-based preferences are somewhat more flexible than proposal-based votes.
3.2 Complementary strategies
Consider an alternative group strategy: the group decides whether to support a proposal using the within-group majority rule. More specifically, the group members vote as individuals; let $n_{G}^{+}({\mathbf{{x}}})$ be the total number of votes they cast for proposal ${\mathbf{{x}}}$ in the internal group voting; finally, each group member $i$ casts $n_{i}^{+}({\mathbf{{x}}})$ votes defined by
$n_{i}^{+}({\mathbf{{x}}})=\begin{cases}1,&n_{G}^{+}({\mathbf{{x}}})>\tfrac{g}{2}\\ \tfrac{1}{2},&n_{G}^{+}({\mathbf{{x}}})=\tfrac{g}{2}\\ 0,&n_{G}^{+}({\mathbf{{x}}})<\tfrac{g}{2}\end{cases}.$ (4)
Such a group can be called majoritarian, as opposed to a utilitarian group. Both group strategies reduce to the individualistic one when applied to a group of one agent.
We say that the strategy of agent $i$ is complementary (or satisfies reversal symmetry) if and only if
$n_{i}^{+}({\mathbf{{x}}})+n_{i}^{+}(-{\mathbf{{x}}})=1\mbox{~{}~{}for any proposal~{}~{}}{\mathbf{{x}}}\in{\mathbb{R}}^{n}.$ (5)
Lemma 2.
The strategies of individuals and members of utilitarian and majoritarian groups are complementary.
Proof of Lemma 2.
Obviously, for a member of a utilitarian group, it holds that $n_{i}^{+}({\mathbf{{x}}})=\tfrac{1}{2}\bigl{(}1+\mathop{\rm sgn}\nolimits\sum_{j=1}^{g}{{x}_{j}}\bigr{)}$, while for a member of a majoritarian group, we have $n_{i}^{+}({\mathbf{{x}}})=\tfrac{1}{2}\bigl{(}1+\mathop{\rm sgn}\nolimits\sum_{j=1}^{g}\mathop{\rm sgn}\nolimits{{x}_{j}}\bigr{)}.$ It is easy to see that in both cases, Eq. (5) is valid. For an individual, this follows as a special case. ∎
3.3 A class of antisymmetric voting bodies
The symmetrized majority rule has a form similar to (4):
$I({\mathbf{{x}}})\;=\;\begin{cases}1,&n^{+}({\mathbf{{x}}})>\tfrac{n}{2}\\ \tfrac{1}{2},&n^{+}({\mathbf{{x}}})=\tfrac{n}{2}\\ 0,&n^{+}({\mathbf{{x}}})<\tfrac{n}{2}\end{cases}.$
This rule satisfies the following condition:
@@ 3.3 A class of antisymmetric voting bodies
Let us say that a stochastic voting rule is complementary iff it satisfies (6).
@@
Any voting body formed by agents with complementary strategies and a complementary voting rule is antisymmetric.
@@
Since agents’ strategies are complementary, for all ${\mathbf{{x}}}\in{\mathbb{R}}^{n}$ we have
$n^{+}({\mathbf{{x}}})+n^{+}(-{\mathbf{{x}}})=\sum_{i=1}^{n}n_{i}^{+}({\mathbf{{x}}})+\sum_{i=1}^{n}n_{i}^{+}(-{\mathbf{{x}}})=\sum_{i=1}^{n}\bigl{(}n_{i}^{+}({\mathbf{{x}}})+n_{i}^{+}(-{\mathbf{{x}}})\bigr{)}=n.$
Since the voting rule is complementary, this implies that $I({\mathbf{{x}}})+I(-{\mathbf{{x}}})=1$ for all ${\mathbf{{x}}}\in{\mathbb{R}}^{n}$. Thus, the voting body is antisymmetric.∎
Figure 2: A non-monotonic complementary voting rule for a society of 6 agents
Another example of a complementary voting rule is shown in Fig. 2.
Lemma 3 can be applied to societies that may contain several groups and employ the symmetrized majority rule.
Corollary 4.
Any voting body formed by individualists, any number of disjoint utilitarian and majoritarian groups, and the symmetrized majority rule is antisymmetric.
The conditions of antisymmetry of a voting body provided by Lemma 3 are not necessary. Indeed, let the agents support all proposals under the voting rule
$I({\mathbf{{x}}})=\begin{cases}\tfrac{1}{2},&n^{+}({\mathbf{{x}}})=n\\ 0,&n^{+}({\mathbf{{x}}})<n\end{cases}.$
Then the voting body is antisymmetric, while neither the agents’ strategy nor the voting rule are complementary. For this voting body, the conclusion of Lemma 1 takes the form ${\mathbb{E}}({\mathbf{D}}_{-{\mathbf{{X}}}})-{\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})=-\tfrac{1}{2}\mu\bm{1}-\tfrac{1}{2}\mu\bm{1}=-\mu\bm{1}$.
3.4 Complementary strategies and rules on opposite proposals
Lemmas 1 and 3 imply
@@ 3.4 Complementary strategies and rules on opposite proposals
For any gain generator ${X}$ with mean $\mu$ and for any voting body formed by agents with complementary strategies and a complementary voting rule, it holds that
@@
The following corollary is a consequence of Theorem 5, Lemma 2 and the fact that the symmetrized majority rule is complementary.
Corollary 6.
The conclusion of Theorem 5 is true for any gain generator ${X}$ with mean $\mu$ and any voting body formed by individualists, any number of disjoint utilitarian and majoritarian groups, and the symmetrized majority rule.
This result underlies the ‘mirrored performance’ effect discussed in Section 4.
@@ 3.5 Gains on location families of generators
Suppose that society consists of agents with complementary strategies and employs a complementary voting rule; let the distribution of the gain generator ${X}^{\mu}$ be symmetric about $\mu.$ Then
@@
Using Lemma 2 we obtain
Corollary 8.
The conclusion of Theorem 7 is true for any gain generator ${X}^{\mu}$ symmetric about $\mu$ and any voting body formed by individualists, any number of disjoint utilitarian and majoritarian groups, and the symmetrized majority rule.
@@ 4 Interpretation and discussion
For any social decision rule $\psi,$ let the expected relative capital gain (ERCG) of an agent under $\psi$ be the mathematical expectation of the difference between the agent’s capital gain under $\psi$ and that under the basic rule after implementing a social decision.
Clearly, ERCG is invariant to replacing rejection with acceptance or any stochastic combination of them at $\mu=0$ in the definition of the basic rule.
For location families (7), consider the ERCG as a function of $\mu.$
@@
For any location family (7) and any agent, the expected capital gain (ECG) under the basic rule is 00 when $\mu\leq 0$ and $\mu$ when $\mu>0.$ Therefore, for any $\mu>0,$ the left-hand side ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}^{-\!\mu}}\bigr{)}$ and the right-hand side ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}^{\mu}}\bigr{)}-\mu\bm{1}$ of (8) are the vectors of agents’ expected relative capital gains (ERCG) under gain generators ${X}^{-\mu}$ and ${X}^{\mu}$ (which are the elements of the location family having means $-\mu$ and $\mu$), respectively. Since by Theorem 7 the left-hand side and right-hand side of (8) are equal, the ERCG of the elements of the family is an even function of $\mu.$
@@
For the society of 21 individualists with the location family of Gaussian generators (7), the expected capital gain under simple majority rule was shown in Fig. 1. The ERCG for this family is presented in Fig. 3a. In addition to a pit of losses in unfavorable environments, this function has a mirror image of that pit in favorable environments. Decisions made by majority votes in these two areas are, on average, disadvantageous to society compared to the basic rule.
The reason for the pit of losses paradox is that in unfavorable environments, the average positive gain is smaller, while the average absolute loss is larger than in the neutral environment of the same variance (where these average values are equal). Therefore, the average difference between [the absolute values of] a gain and a loss is negative. Consequently, the proposals accepted by simple majority with a small margin of votes in favor (which is typical in an unfavorable environment) yield, on average, a total loss. Thus, in environments with $\mu<\mu_{0}<0,$ where $\mu_{0}$ is some threshold, the majority’s gain does not compensate the loss suffered by the minority. Note that this imbalance identified within the ViSE framework is relevant to practice.
@@
Thus, if for the society with parameters corresponding to Fig. 3a, environments with $\mu/\sigma<-0.254$ are considered ‘hell’ and environments with $\mu/\sigma>0.254$ are considered ‘heaven’, then simple majority rule is worse than the basic rule in both of these realms. Simple majority rule outperforms the basic rule only in a neighborhood of the neutral environment ($\mu=0$), where the ERCG peak is quite high.
Let us consider another society: with a utilitarian group of 7 agents and 14 individualists (Fig. 3b). For it, Gaussian gain generators produce a completely positive ERCG for group members (blue curve). In contrast, the ERCG of individualists is negative, except for a narrow interval of $\mu/\sigma\in[-0.03,0.03]$ (red curve).
It is instructive to compare this with the ERCG for the societies of 21 individualists (red dotted curve) and of 21 group members (blue dotted curve). The most advantageous position is in the group of 7 agents and the worst is among the individualists complementing this group. Such a group can be considered an ‘elite’. It usually receives relatively high capital gains at the expense of individualists. It should be noted that the ERCG of the whole mixed society (gray curve) is even lower than that for the society of 21 individualists. The impact of a ‘responsible elite’ on public welfare was investigated by [43, 44]. In most cases, such an elite provides society with capital growth.
Obviously, the most beneficial strategy for the whole society is to be one cohesive group. However, in practice this is unrealistic. While a group of like-minded people is able to develop a common objective function, this is problematic for society. Therefore, non-consensus-based decisions are of primary interest. The ‘basic rule’, which outperforms the simple majority in both ‘hell’ and ‘heaven’ is extremely crude. How can taking into account voters’ opinions improve the outcome? Such a social decision rule is voting with the optimal threshold ([10, 26]). This threshold is exactly $0.5$ in the neutral environment, as well as for $\mu/\sigma\in[-0.1,0.1]$ in our example (Fig. 3a). Outside a certain interval around 00, the optimal threshold is usually (but not always) higher in unfavorable and lower in favorable environments. The ERCG curve in the case of voting with the optimal threshold clearly has no pit of losses. For the society of 21 individualists, it is shown in Fig. 3a.
@@ 5 Conclusion
The ViSE model is designed to evaluate the effectiveness of individual voting strategies and social decision rules in environments of varying favorability and, more generally, in environments with various gain distributions. In this paper, we show that in the case of Gaussian environments and two-component (a group and individualists) societies, the symmetrized majority rule is effective only under more or less neutral conditions, while in ‘heaven’ or ‘hell’ it is harmful. Furthermore, for environments with symmetric distributions, the performance of this rule is symmetric about the neutrality point. For general distributions with finite mean, the performance is the same for any two opposite generators. This is true both for the class of voting bodies formed by agents with complementary strategies and a complementary voting rule, and for the even more general class of antisymmetric voting bodies. Returning to two-component societies: in some of them, cooperating group members improve their position at the expense of individualists, which is disadvantageous for the whole society.
@@
Acknowledgments We are very grateful to two anonymous referees for their careful reading and especially for suggesting a generalization and the idea of reverse time to obtain a simpler proof of the technical result. In the final version, we use a similar approach in other terms, supplemented by the condition of antisymmetry of the voting body, which is necessary to generalize the result. We also thank the participants of the seminar on stochastic voting models at the Institute of Control Sciences of the Russian Academy of Sciences: Yana Tsodikova, Anton Loginov, Vitaly Malyshev, Vladislav Maksimov, Anna Khmelnitskaya, Irina Kalushina, Zoya Lezina, Vladimir Borzenko.
@@
Funding   Work of P.C. was supported by the Israel Science Foundation (grant No. 1225/20) and European Union (ERC, GENERALIZATION, 101039692). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.